-4x - 5y - 8z = 268
-10x + 7y + 6z = -786
-12x + 9y - 2z = -548
I could really use some help on this problem. Thanks!!

Answer by stanbon(75887)

-4x - 5y - 8z = 268
-10x + 7y + 6z = -786
-12x + 9y - 2z = -548
---
Problems like this are a waste of a lot of time.
--
Hopefully you have a calculator to work this
is 1/100 th the time it takes to do them by hand.
----
x = 41
y = -16
z = -44
=================
Cheers,
Stan H.
===============